Calculate the price elasticity of demand for the following products and state whether demand is price elastic, inelastic, or unit elastic,
a. Raw sugar prices rose by 3% and raw sugar consumption declined by 2%.

b. A 10% increase in the price of cigarettes reduce smoking among 8th to 10th graders by 10%.

c. The price of luxury cars increases by 4% while the quantity demanded decreases by 8%.
Answer #1

a) e= (% change in quantity demanded)/(% change in price) = -2/3 = -0.67, Since the absolute value of elasticity of demand is 0.67 which is less than 1, the demand is inelastic, i.e. for every percentage increase in price, the demand falls by less than 1%.

b) e= (% change in quantity demanded)/(% change in price) = -10/10 = -1, Since the absolute value of elasticity of demand is 1, the demand is unit elastic, i.e. for every percentage increase in price, the demand falls by 1%.

b) e= (% change in quantity demanded)/(% change in price) = -8/4 = -2, Since the absolute value of elasticity of demand is 2, the demand is elastic, i.e. for every percentage increase in price, the demand falls by 2%, i.e. demand falls by twice the increase in prce.
Answer #2

a)price elasticity = -(dQ/Q)/(dP/P) = -2/3 = inelastic

b) = -10/10 = -1 unitary elastic

c) -8/4 = -2 = elastic